A body is dropped from height of 100m at what height the velocity of body will be equal to one half of velocity when it hits the ground

Grade:12th pass

1 Answers

Lohith
6 years ago
At a height if h/4 the velocity will be half of the velocityProof V = sqrt of 2gh = sqrt(2*10*100) = 10*sqrt (20) V/2 = 5*sqrt(20) --------- (1)Consider v2 - u2 = 2ghTherefore (v/2)2 - 0 = 2gh (V/2)2 = 2*10h(5*sqrt (20))2 = (2*10*h)25*20 = 20*hThere fore h = 25m